Institut for Matematik og Datalogi Odense Universitet February 14, 1995

JFB

DM15 – F95 – Week 17

Lecture, February 6

We covered section 8.3 and through example 8.4.1 of section 8.4 in Sudkamp.

Lecture, February 13

We will finish chapter 8 and begin chapter 9 in Sudkamp, covering section 9.1 and possibly 9.2.

Lecture, February 20

We will finish most of chapter 9 in Sudkamp.

Problems to be discussed in week 8

• Do problem 16 on page 208 of Sudkamp.

• Do problem 18 on page 209 of Sudkamp.

• Do problem 23 on page 209 of Sudkamp.

• Do problem 2 on page 232 of Sudkamp.

• Do problem 4 on page 232 of Sudkamp.

• Do problems 6c and 6d on page 233 of Sudkamp.

A problem for which you can get comments to your solution

• Do problem 8 on page 233 of Sudkamp. (Your answer to this problem can be turned into your instructor, Peter Høyer, in week 8.)